### v1

Abstract
In this second paper on the variability survey of central stars of planetary nebulae (CSPNe) using ZTF, we focus on the 11 long-timescale variables with variability timescales ranging from months to years. We also present preliminary analyses based on spectroscopic and/or photometric follow-up observations for six of them. Among them is NGC 6833, which shows a day periodic variability with strange characteristics: ‘triangle-shaped’ brightening in , , and WISE bands but almost coincidental shallow dips in the -band. We speculate this to be a wide but eccentric binary with the same orbital period. Long-period near-sinusoidal variability was detected in two other systems, NGC 6905 and Kn 26, with periods of days and days, respectively, making them additional wide-binary candidates. The latter also shows a short period at hours which can either be from a close inner binary or pulsational origin. We present CTSS 2 and PN K 3-5 which show brightening and significant reddening over the whole ZTF baseline. A stellar model fit to the optical spectrum of CTSS 2 reveals it to be one of the youngest post-AGB CSPN known. Both show high-density emission-line cores. These appear to be late thermal pulse candidates, currently evolving towards the AGB phase, though alternative explanations are possible. We then present recent HST/COS ultraviolet spectroscopy of the known wide-binary candidate LoTr 1 showing that the hot star is a spectroscopic twin of the extremely hot white dwarf in UCAC2 46706450. We think that the long photometric period of 11 years is the binary orbital period. Finally, we briefly discuss the ZTF light curves of the remaining variables, namely Tan 2, K 3-20, WHTZ 3, Kn J1857+3931, and IPHAS J1927+0814. With these examples, we present the effectiveness of the von Neumann statistics and Pearson Skew-based metric space in searching for long-timescale variables.
1 Introduction
Over the past decade, studies of photometric variability of central stars of planetary nebulae (CSPNe) have significantly improved our understanding of these systems. The majority of past works have been dedicated to the search for binarity in CSPNe in support of the common-envelope (CE) formation channel of planetary nebulae (PNe; see, for example, Bond et al. 1978 and Bond & Livio 1990 for some of the first discoveries). Wide-field photometric surveys like the Optical Gravitational Lensing Experiment (OGLE), Kepler-K2, Transiting Exoplanet Survey Satellite (TESS), and recently the Zwicky Transient Facility (ZTF) have together discovered more than a hundred binary CSPNe (Miszalski et al., 2009a, b; Jacoby et al., 2021; Aller et al., 2020, 2024; Bhattacharjee et al., 2025; Chen et al., 2025). Periodicity has also led to the characterization of other phenomena like rotational variability caused by stellar spots (see for example Werner et al. 2019, 2023; Bond & Zeimann 2024), or pulsations of the hot CSPN (Bond & Meakes, 1990; Ciardullo & Bond, 1996; Hajduk et al., 2014; Córsico et al., 2021).
Though much less explored compared to periodic variables, CSPNe also exhibit aperiodic variability. One of the causes is strong winds in CSPNe, mostly causing ‘jittering’ variability (see for example Handler et al. 1997; Arkhipova et al. 2012, 2013; Córsico et al. 2021). Unusual and large-amplitude aperiodic photometric variability of CSPNe has also led to the discoveries of R Coronae Borealis variables inside PNe, providing significant insights into the late thermal pulse formation channel of this class of stars (Gonzalez et al., 1998; Asplund et al., 1997; Clayton & De Marco, 1997; Jeffery & Hambsch, 2019; Rosenbush & Efimov, 2015). Recently, transits of dust and debris discs was proposed as an explanation for the deep quasi-periodic dips seen in the nucleus of the planetary nebula WeSb 1 (Bhattacharjee et al., 2025; Budaj et al., 2025). These demonstrate the variety of causes that can induce photometric variability in CSPNe.
Most of the aforementioned phenomena occur on short timescales, ranging from a few minutes to a few days. On the other hand, there have only been very few instances of long-timescale photometric modulation in CSPNe. One example is PN IC 4997 (Aller & Liller, 1966) which shows decade-long brightening and dimming periods which may result from either a long-period binary (Kostyakova & Arkhipova, 2009) or wind/mass loss variations (Arkhipova et al., 2020). Long-period light curve modulation due to binarity is seen in LoTr 5 (period of 7.4 years, Kővári et al. 2019). Long-term photometric monitoring also resulted in the discovery of the day binary in pre-PN V510 Pup (Manick et al., 2021). Post-asymptotic giant branch (AGB) stars undergoing (very) late thermal pulses also show long-timescale photometric changes (see for example van Genderen & Gautschy 1995 for FG Sge and Schaefer & Edwards 2015 for Stingray Nebula). There are a few other similar examples in the literature. However, there has not been any systematic search for long-timescale variability in CSPNe.
This is the second of a two-paper series conducting a systematic study of CSPN variability with Zwicky Transient Facility (ZTF, Graham et al. 2019; Masci et al. 2019). Here, we focus on the CSPNe showing long-timescale photometric variability in ZTF (the first one focused on the short-timescale variables, Bhattacharjee et al. 2025, hereafter Paper I). We find 11 objects showing prominent long-term variations in the light curve, making this by far the largest study of long-timescale variability in CSPNe. The rest of the paper is as follows. In section 2, we briefly review the methodology employed in selecting the variable sources. In section 3 we present our results. This includes spectroscopic observations of some of the objects and related analyses. Finally, we present our conclusions in section 4.
2 Method and Sample Selection
The detailed procedure for the selection of variable CSPNe in ZTF data has been described in Paper I. We briefly review the methodology here. Our starting sample is the Gaia-EDR3-selected list of CSPNe presented in Chornay & Walton (2021). We restrict ourselves to sources with CSPN identification reliability , resulting in 1812 CSPNe, of which ZTF data is recovered for 991 sources. Following this, we applied several photometric quality cuts, primarily to avoid sources where the extended nebula significantly affects the photometry. This reduced the number of sources to 490. To quantify variability, we calculate the Normalized Excess Variance metric (NEV) for the light curves. Using an appropriate cut on this metric value, we arrive at our final list of 94 highly variable CSPNe (HNEV sample). Among the HNEV sources, we classify 83 objects to be short-timescale variables (the focus of Paper I). The variability of the remaining 11 objects occurs on much longer timescales (months to years) and is well resolved with ZTF cadence. We classify these objects as long-timescale variables, which are the focus of this paper. The list of the 11 long-timescale variables is presented in Table 3.
In Paper I, we had presented a two-dimensional metric space defined by the metrics von Neumann statistic (vonN) and Pearson Skew (SkewP) as a simple and novel technique to identify “exotic variables”. Briefly, the former metric quantifies the orderliness in the variability (against random variability from photon noise), and the latter differentiates between dimming and brightening events. We discussed that objects with low vonN () are often “interesting”, i.e., objects which can teach us new phenomena. We demonstrated that the debris disk candidate WeSb 1 lies in the low vonN space and its large negative SkewP successfully represents the deep eclipses. In Appendix A, we identify the long-timescale variables in this space and show its effectiveness in identifying them from large data sets. Such a discovery tool is expected to come in handy with upcoming large datasets with the advent of new surveys like Large Synoptic Survey Telescope (LSST, Ivezić et al. 2019).
3 Results and Analyses
PNG | Name | RA | Dec | GaiaID | PNStat | PNRad | Rel | Known Period | ZTF Period |
082.5+11.3 | NGC 6833 | 297.444 | 48.961 | 2086764737566091520 | T | 0.3 | 1.0 | – | 978 d |
061.4-09.5 | NGC 6905 | 305.596 | 20.104 | 1816547660416811008 | T | 21.65 | 1.0 | – | 700 d? |
084.6-07.9 | Kn 26 | 320.789 | 38.97 | 1965268354805776896 | T | 55.0 | 1.0 | – | 230 d? 1.18 hr |
044.1+05.8 | CTSS 2 | 282.695 | 12.625 | 4504643171308707072 | T | – | 0.99 | – | – |
034.3+06.2 | K 3-5 | 277.941 | 4.086 | 4283434037501296128 | T | 5.0 | 0.99 | – | – |
228.2-22.1 | LoTr 1 | 88.778 | -22.901 | 2917223705359238144 | T | 71.0 | 0.78 | 11.28 yr (1), 6.48 d (2) | 6.48 d |
069.6+15.7 | Kn J1857.7+3931 | 284.426 | 39.517 | 2103243049609815552 | P | 75.0 | 1.0 | – | – |
032.5-03.2 | K 3-20 | 285.542 | -1.813 | 4262011019866089984 | T | – | 0.98 | – | – |
014.0-02.5 | Tan 2 | 276.321 | -17.941 | 4096278711670069248 | T | 3.0 | 0.83 | – | – |
045.0-12.4 | WHTZ 3 | 299.497 | 4.792 | 4248349003725092864 | T | 46.0 | 0.99 | – | – |
044.3-04.1 | IPHAS J192717.94+081429.4 | 291.825 | 8.241 | 4296182638654047232 | P | – | 1.0 | – | – |
Name | Summary |
NGC 6833 | ‘Triangle-shaped’ periodic brightening in ZTF , and WISE bands associated with shallow dips in . Eccentric binary with periastron interaction? |
NGC 6905 | Long periodic ( days, also days?) near-sinusoidal variability. Wide binary? |
Kn 26 | Long periodic ( days, also days?) near-sinusoidal variability: wide binary? Short period ( hours) arising from either pulsation or close binary. |
CTSS 2 | Steady brightening associated with significant reddening. Very young pre-white dwarf CSPN ( kK and ). Late thermal pulse? |
K 3-5 | Very similar light curve behavior as CTSS 2, but no information about the CSPN. Late thermal pulse? Winds and dust? |
LoTr 1 | A very hot ( kK) white dwarf with a rapidly rotating cool giant ( and K) companion. Orbital period of 11 years? |
Kn J1857+3931 | Long-timescale dimming over the whole ZTF baseline. |
K 3-20 | Long-timescale brightening (symmetric in appearence) associated with redenning. Part of a longer-periodic binary light curve? |
Tan 2 | Pointed dip in the beginning of ZTF light curve. Undergoing dip again after years. Wide-orbit debris eclipses or binary? |
WHTZ 3 | Short amplitude likely periodic long-term modulation in ZTF. Wide binary? Noisy data prevents detailed analyses. |
IPHAS J1927+0814 | Erratic high amplitude variability. Reason unclear: dust and debris? accretion? |
Below we present some discussion on the long-timescale variables. The wide variety of light curves (and, thus, the underlying causes of variability) makes the detailed study of all the objects together difficult. We are in the process of obtaining follow-up spectroscopic and photometric observations of these objects. As a demonstration of this effort, we present the preliminary analyses and discussion on several of the objects based on the available data. Before diving into the details, we summarize our analyses in Table 2 for the convenience of the readers.
We note here that except WHTZ 3, the ZTF variability of all the other objects below were recently reported in a completely independent work by (Chen et al., 2025, henceforth, C25). Though they successfully present a broad range of discoveries from ZTF with much less stringent cuts than the ones applied here (going beyond the HNEV sample in our papers), they provide no additional discussion on the discovered systems. This makes this work different as we present follow-up observations and extensive discussions. These shall serve as the first step in understanding these complex objects. But we refer to C25 whenever needed in our discussions.
All of these objects appear in the catalog of PN (candidates) detected by GALEX (Gómez-Muñoz et al., 2023). Five among these objects (LoTr 1, Kn J1857+3931, NGC 6833, NGC 6905, and Kn 26) also appear in the catalog of hot subluminous stars by Culpan et al. (2022).
3.1 NGC 6833
NGC 6833 is a well-known PN and a True PN in HASH222Hong Kong/Australian Astronomical Observatory/Strasbourg Observatory H Planetary Nebula (HASH, Parker et al. 2016) catalog is the largest catalog of PNe which serves as an important reference for Paper I and this work. It ranks its PN candidate entries as “True”, “Likely” and “Probable” PNe, in the decreasing order of their likeliness of actually being a PN. The True PNe are spectroscopically verified to be so. with an apparent diameter of . Due to its compact angular size, current ground-based surveys cannot resolve the central star from the nebula. Thus, the recorded fluxes are a combination of both the star and the nebula. Images from the Hubble Space Telescope (see for example Hyung et al. 2010) reveal the box-like shape of the PN. NGC 6833 has found its presence in various catalogs over decades and has been the subject of several spectroscopic studies. The photometric behavior, however, remained unnoticed.
3.1.1 Light curves
The ZTF light curve for NGC 6833 is shown in Figure 1 (top panel). Fortunately, this target has good -band coverage. The most prominent feature is the regular brightening events in the ZTF and the redder bands, with symmetric ingress and egress. The amplitude is in the range of , with the largest amplitude in -band. We note here though that the nebular flux can reduce the perceived variability, so the intrinsic amplitude of the source (likely the central star) can be much higher (see a brief discussion in Section 3.1.3). The less prominent but more intriguing feature appears in the ZTF band, which exhibits shallow (amplitude of ) dips nearly coincident with the peaks in the redder bands. The behavior is periodic and significant peaks appear in the Lomb-Scargle periodogram in all the three ZTF bands. The result is shown in the top panel of Figure 2. The detected periods in the three bands are marginally different (which is not surprising with long periods), with a mean value of days. The phase-folded ZTF light curve at this mean period is shown in the middle panel of the same figure. Here, the activity appears to be slightly phase-delayed from the redder to the bluer bands. However, the current data is not enough to establish this property with high confidence.
We also present the WISE light curve in Figure 1. For better presentation, we only show the neoWISE light curve in this Figure. The periodic brightening events are evident and coincident with the activity in ZTF. Unlike ZTF, the variability amplitude in both W1 and W2 appear similar. The lomb-Scargle periodogram applied to the WISE data shows a similar period of days, as in ZTF. For consistency, we use the ZTF-inferred mean period to phase-fold the WISE data. The result is shown in the bottom panel of Figure 2. Except for the first data point (which happens to be the first allWISE reading and likely outlier affected by systematics relative to neoWISE), the phase folded light curves closely resemble that of the red bands in ZTF.
Before any further discussion, we caution that the brightness of this object ( mag in and mag in ) is very close to the saturation limit of ZTF ( in all bands). In fact, in one of the ZTF fields, the reference image is indeed saturated in the -band (where the object is brightest, mainly due to the very strong [O III] emission lines). Though such data points get rejected through the imposed quality cuts, the chance remains that the other flux readings are unreliable. To ascertain that this has not significantly affected the observed variability in the -band, especially the dip, we checked the -band light curve from the All Sky Automated Survey for SuperNovae (ASAS-SN) light curve provided at Sky Patrol V2.0333http://asas-sn.ifa.hawaii.edu/skypatrol/ (Shappee et al. 2014; Hart et al. 2023), which has a larger saturation brightness ( magnitude). The ASAS-SN -band light curve displays similar behavior as ZTF and the dips are visible. The data quality, however, is worse than ZTF thus not shown. For ZTF and bands, manual inspection shows that neither the reference nor the science images are saturated. We thus proceed with the assumption that the light curves are trustworthy.
3.1.2 Follow-up Optical Spectroscopy
NGC 6833 was observed with the DouBle SPectrograph (DBSP; Oke & Gunn 1982) attached to the Cassegrain focus of the Palomar 200-in Hale Telescope. We used the D55 dichroic, the 600 line mm-1 grating for the blue arm blazed at 3780 Å, and the 316 line mm-1 grating for the red arm blazed at 7150 Å. We used grating angles of 27∘17′ and 24∘38′ for the blue and red sides, respectively. With these setups and a slit width of 1.5′′ long-slit, we achieve resolving powers of in the blue arm and in the red arm. These setups provide continuous spectral coverage across 2900–10800 Å, with the division between blue and red arms typically occurring around 5650 Å. We followed the same reduction procedure as described in Section 4.2.1 in Paper I. Briefly, we used PypeIt (Prochaska et al., 2020a, b) with the options use_2dmodel_mask = False and no_local_sky = True to avoid masking and local subtraction of the strong emission lines. Following this, we perform manual trace identification on the two-dimensional spectrum. Figure 3 shows the optical spectrum for NGC 6833.
As mentioned previously, the object is spectroscopically well-studied (for example, Wright et al. 2005; Wesson et al. 2005; Hyung et al. 2010). However, given the knowledge of the photometric variability, the epoch of this spectrum is “special” as the object is nearly at its photometric extremum. We attempted to estimate the phase at which the previous spectroscopic observations were made, using the observation dates mentioned in the corresponding papers. Assuming that the ZTF inferred period is stable, the best estimate places the optical observations of Wright et al. (2005) and Hyung et al. (2010) at the quiescent phase. However, no significant difference in the spectral features was observed compared to the earlier studies. The calculated emission line fluxes of the major emission lines are also in very good agreement with those previously reported (in Wright et al. 2005, for example). However, a more thorough comparison is needed to identify differences at the level that can induce the observed variability (which includes ensuring consistency of trace selection, seeing, instrumental effects, etc.). This is beyond the scope of this work.
3.1.3 Discussion
The stable variability and the associated periodicity strongly suggest binarity, with the period of variability as the binary orbital period. Long-period binaries inside PNe are only rarely observed, making NGC 6833 an important addition. Only three CSPNe are known to have an orbital period longer than the inferred period of NGC 6833: BD+332642 (Van Winckel et al., 2014; Jones et al., 2017), LoTr 5, and NGC 1514 (Jones et al., 2017). The shape of the light curve and the larger amplitudes in the redder bands suggest an irradiation effect. In that case, the relatively high variability amplitude (at this period) requires the orbit to be eccentric, which is not unexpected in long-period binaries. All the other known long-period binaries in CSPNe are significantly eccentric. Eccentricity pumping mechanisms like orbital phase-dependent mass transfer can indeed maintain high eccentricities up to periods of days (see for example Bonačić Marinović et al. 2008). At this period, it is also at a margin of being a post-CE system.
Even within the binary scenario, mechanisms other than irradiation effect can possibly contribute to the observed variability. For example, the light curves in ZTF- and redder bands bear a strong resemblance to several RV Tauri (evolved pulsating giants) systems showing long-term photometric modulation (known as the RVb phenomenon, see for example the light curves in Kiss & Bódi 2017 which shows very similar ‘triangle-shaped’ brightenings with similar periods in the range of a few years). The widely accepted explanation is that the variability results from the phase-dependent dust-obscuration of the flux-dominant star in the presence of a thick circumbinary disc (Van Winckel et al., 1999; Gezer et al., 2015). In this case, too, the photometric period represents the orbital period. RVb systems, in fact, are often regarded as CE progenitors (see for example Kiss & Bódi 2017). Binaries with similar periods and eccentricity have also been detected in post-AGB systems (van Winckel et al., 2009; Oomen et al., 2018).
However, the above scenarios individually fail to account for the anti-correlated behavior in the and the redder bands. A combination of both effects, however, could explain the variability. In an eccentric orbit, interaction (either directly between the stars or with a circumstellar/binary disc) can be enhanced near periastron. This can lead to the significant production of dust, leading to increased extinction in the blue bands. In the red bands, both irradiation of the cooler component and enhanced dust emission can contribute to the brightening. The stochastic nature of the interaction may also explain any offset between the peaks in the red and the dip in the blue. Note that anticorrelation between ZTF- and bands were also recently reported in another system (PM 1-322, Paunzen et al. 2023, now classified to be a symbiotic candidate on HASH), and a similar scenario was proposed as a possibility.
We now briefly consider possible cases where the photometric period does not relate to binarity. Rotational modulation is often a cause for periodic variability, but, in most cases, the light curves in different bands are in phase. C25 compares it with light curves of a group of magnetic and chemically peculiar stars called Canum Venaticorum (ACV), which do indeed show anti-correlation between blue and red bands (see for example Gröbel et al. 2017; Faltová et al. 2021). However, the periods of these stars are much shorter (all known objects having days, Faltová et al. 2021) compared to that inferred in NGC 6833. On a different possibility, broadly similar light curves are witnessed in systems involving Be stars losing mass periodically due to pulsational instabilities. The lost mass appears as cold dust emitting mostly in far optical to mid-IR, leading to the brightening. Thus, it might be possible for the dust to induce a dip in the blue bands through extinction. In most such cases however, the ingress time is significantly shorter than the egress time and the infrared amplitudes are much larger than observed in NGC 6833 (see, for example, the light curves presented in Froebrich et al. 2023). This makes this scenario, too, unlikely.
All the above scenarios are expected to result mostly in the variability of the stellar continuum. However, any variation of the very strong emission lines can also induce significant photometric variability. We use the spectrum to check if the observed variability may result solely from emission line variations. We consider the ZTF- band () which shows the largest amplitude of . We calculate the emission line fluxes by appropriately modeling and subtracting the continuum. The emission line (including the Paschen series) contribution to the total flux is . With this spectrum taken at the -band maximum, this would require a significant decrement in the line strengths of all the emission lines to contribute meaningfully to the variability, which is unlikely. This, however, may not hold for the other bands, especially where the emissions are strong. In the ZTF- and bands, for example, the emission line comprises and of the total flux. Additionally, a significant fraction of the continuum is expected to be of nebular origin. Further epochal spectroscopic observations and proper modeling of the nebular continuum are needed to identify the source of the variability.
We briefly mention here that, in case the photometric variation is of stellar origin, the strong emission lines (and possibly the nebular continuum) can significantly suppress the true variability of the CSPN. We test this by assuming that the variability arises solely from the continuum. We calibrate the ZTF flux to the DBSP spectrum (in the respective bands) followed by subtraction of the emission line fluxes to get revised variability amplitudes. The amplitude of the dip in the increased to . The peaks in the and bands increased to relative flux values of and respectively. Thus, the band may not have an intrinsically higher amplitude than the band, as appears in the raw light curve. Though this exercise is preliminary and can only yield crude estimates, it shows that the effect can be significant. Interpretations, thus, need to be made with sufficient care.
3.2 NGC 6905 and Kn 26
NGC 6905 is a spectroscopically well-studied planetary nebula. The central star of NGC 6905, HD 193949, is a [WR]-type star with a temperature of kK (Gómez-González et al., 2022). The spectral classification seems to be uncertain, with reports of it being both [WO] (for example see Gómez-González et al. 2022) and [WC] (for example see Córsico et al. 2021) existing in the literature. The star also shows pulsations with periods in the range of seconds (Ciardullo & Bond, 1996; Córsico et al., 2021). The nebula shows a bipolar morphology. Kn 26, on the other hand, is a known quadrupolar PN (Guerrero et al., 2013), with Lanning 384 being the central star. Guerrero et al. (2013) finds the blue-side SED to be consistent with a blackbody of kK. Unlike the other, Kn 26 has not previously been studied spectroscopically.
3.2.1 Light Curves
Both NGC 6905 and Kn 26 show long-term modulation in the ZTF light curve. The Lomb-Scargle periodogram over the long-period range (same as mentioned in caption of Figure 2) shows a very prominent peak at days and a weaker peak at days for NGC 6905. C25 reports the longer period in their work. For Kn 26, two peaks of similar powers at days and days (the former being slightly dominant) appear in the periodogram444C25 does not report the long-period for Kn 26. One might guess that this is due to the use of standard photometry against difference photometry. However, the long periods in both the objects are also recovered from the standard photometry data. This in turn shows that the photometric behavior is not an artifact of difference imaging.. The periods appear individually in the and bands (see Figures 4 and 5, top panel). It is interesting to note that the inferred periods in the two objects are quite similar. This, and the closeness of the periods to possible systematic timescales (seasonal breaks and yearly observation cycles) prompted a few tests on the reality of the variability/periodicity. The details of the tests (which includes investigating the ATLAS data) can be found in Appendix B. We find no reason to reject the periods. Continued photometric monitoring is needed to confirm both the periodicity and the significance of the two peaks in the periodogram.
In addition to the long period, Kn 26 also shows a short-period signal at hours. When a Lomb-Scargle analysis over the short-period range (200 seconds to 100 days, same as used in Paper I) is applied on the raw light curve, the corresponding peak is visible, but it is subdominant to the -day cadence peak. However, the peak appears dominantly when the long-period modulation is subtracted out from the light curve and the periodogram is re-run. This result is presented in the third panel in Figure 5. The period is independently inferred in both the ZTF bands.
We used follow-up photometric observations to verify this period. Kn 26 was observed with the Wide Field Camera (WFC) at the Isaac Newton Telescope (INT) in July, 2021 in the -band with s exposure, and in August, 2022 in - and - bands with s and s exposures, respectively. The data were debiased and flat-fielded, before performing differential photometry of the central stars against non-variable field stars, all using standard astropy routines (Astropy Collaboration et al., 2018). Recently, we re-observed Kn 26 with the Caltech HIgh-speed Multi-color camERA (CHIMERA) in and bands on September, 2024. The CHIMERA images were corrected using bias and flat images555https://github.com/caltech-chimera/PyChimera. Photometry was extracted using aperture photometry (with a variable aperture size) using a modified version of the ULTRACAM pipeline (Feline et al., 2005). We made a differential light curve using a single reference star. GPS timestamps were used to determine the mid-exposure time of each image.
A Lomb-Scargle periodogram applied to the INT band data yielded the same period as ZTF (fourth panel in the same figure). The morphology of the phase-folded light curve also resembles that of ZTF. Owing to fewer (and denser) observations with INT and bands and CHIMERA, we do not apply the period search. Rather, we show the full light curves in the bottom-most panel of the same figure. The INT light curves show a flat top with pointed periodic dips, in agreement with the INT -band and ZTF light curve shapes. But the CHIMERA light curves appear markedly different. However, we note that there were some issues with the instrument which includes it going out of focus for a while (the shaded portion), which may have degraded the quality of the data. The periodicity, however, is visually evident in the light curves.
3.2.2 Follow-up Optical Spectroscopy
As noted earlier, the CSPN of NGC 6905 is spectroscopically very well studied. This is not the case with Kn 26, which prompted us to perform a spectroscopic observation. We took four 900 s exposures of Kn 26 at the twin 8.4m Large Binocular Telescope (LBT) using the Multi-Object Double Spectrographs (MODS, Pogge et al. 2010) on December, 2024. MODS provides two-channel grating spectroscopy by using a dichroic that splits the light at Å into separately optimized red and blue channels. The spectra cover the wavelength region Å and Å with a resolving power of and , respectively. The spectra were reduced using the modsccdred666https://github.com/rwpogge/modsCCDRed PYTHON package (Pogge, 2019) for basic 2d CCD reductions, and the modsidl777https://github.com/rwpogge/modsIDL pipeline (Croxall & Pogge, 2019) to extract 1d spectra and apply wavelength and flux calibrations.
The coadded MODS spectrum is shown in Fig. 6. We note that the Balmer and He II lines appear to be a super composition of a narrow emission line (likely resulting from the PN) and weaker but broader emission line (possibly originating from the photosphere). We identify the N III C III C IV emission line complex around 4650 Å but note that the the C IV 5801, 5812 Å doublet is not visible. Therefore, Kn 26 can be considered as a weak emission-line type ([WELS]) central star (Marcolino & de Araújo, 2003). We note here that the emission lines can also arise from the irradiated companion (as often the case with WELS type PN nuclei, see for example Miszalski et al. 2011a), and can be an indirect confirmation of the binarity of the system.
3.2.3 Discussion
The possible reasons behind the long-period variability observed in both the objects are, at this stage, uncertain. Binarity often results in long-term photometric modulation though ellipsoidal variability and/or irradiation effect. The spectrum of Kn 26, for example, might speak in favor of the latter as the weak emission lines may originate from an irradiated companion. Other causes of photometric variability is also possible, for example through stellar surface activity phased with the orbit, as suspected for the wide-binary LoTr 5 (see Kővári et al. 2019). In all of these cases, the photometric period is directly related to the orbital period (equal or double). Thus, both NGC 6905 and Kn 26 stand as candidates for long-period (maybe post-CE) wide binary systems. On a different possibility, long-timescale quasi-periodic modulation may result from close-binary involving a magnetically active giant companion: RS CVn systems. However, the associated long-periods are usually much longer ( years, see Martínez et al. 2022) than seen in NGC 6905 and Kn 26.
The light curves also bear a resemblance to another class of phenomena: Long Secondary Period (LSP) photometric modulations observed in several RGB and post-AGB stars (for example, see Kiss et al. 2006, Nicholls et al. 2009, and Soszyński et al. 2021). The phenomenon is not yet well understood and several possibilities have been proposed over the past couple of decades. One is related to binarity, where the companion is a low-mass object (large planet or brown dwarf), which carries a dust cloud around it, causing the occultation of the star (Soszyński et al., 2021). In such a scenario, we would expect an infrared excess in the spectral energy distribution (SED) of the CSPN. Interestingly, strong infrared excess is seen in both NGC 6905 and Kn 26 (Figure 7). However, this excess is unlikely to be caused by a localized dust cloud around a companion, as the dimensions of the cloud required to match the observed excess are large. For example, the mid- and far-infrared excess in NHC 6905 is matched with a blackbody of temperature of K and radius of AU. This still does not rule out this possibility as the dust cloud causing the photometric variation might just be a smaller embedded component.
Another explanation for LSPs is pulsations of RGB and AGB stars (see for example Wood et al. 2004). It is not straightforward to extend this interpretation to CSPNe, as the white dwarfs are too compact to oscillate at such long periods. Their oscillation periods are usually much shorter (on the order of minutes, as in case of NGC 6905). This may again indicate binarity with a pulsating RGB and AGB companion. This is unlikely but definitely still a possibility.
Pertaining to the short period, an explanation can be a close binary. The phase folded light curve is characterized by “pointed” minimas, morphologically similar to ellipsoidal variability with grazing eclipses. In that case, the orbital period is 2.36 hr, still making it the candidate for the shortest known binary period. In that case it would either be a double degenerate or a white dwarf – subdwarf system. Assuming typical masses of these stars yields a orbital semi-major axis of (consistent with ellipsoidal variability) and velocity (of each star) to be more than . This should be detectable if the inclination is sufficiently high. However, given that we do not see deep eclipses might indicate a low inclination, and thus, a low percieved radial velocity. The LBT spectra at different epochs do not show any obvious radial velocity shift. Thus this scenario too remains to be confirmed. But, if the short period indeed signifies a binarity, the interpretation of the long period becomes even more interesting and might indicate the possibility of a triple system inside a planetary nebula.
3.3 CTSS 2 and PN K 3-5
CTSS 2 (PN G044.1+05.8) is a True PN on HASH. The nebula is faint and can be seen only in deep narrow-band images (see Figure 12.8 in Stanghellini et al. 2016 for the HST image). There has been some uncertainty in the literature regarding the true nature of the central star of CTSS 2. Stanghellini et al. (2016) notes that the identified CSPN appears too bright to be associated with the faint and evolved PN, and considers the possibility of it being an unrelated field star. They also consider the possibility of it being a binary, as does Moreno-Ibáñez et al. (2016). On the other hand, García-Hernández & Górny (2014) found emission signatures unusual for a PN in the Spitzer IR spectrum, and identified CTSS 2 to be a symbiotic system candidate. Confirmatory signatures of it being either a PN or symbiotic system are, however, lacking.
PN K 3-5 (PN G034.3+06.2) is also a True PNe on HASH. It is a resolved bipolar nebula, with an angular diameter of . The most recent spectroscopic study was performed in the spectral survey of central stars by Weidmann et al. (2018). They indicated the presence of emission lines both from the central stellar region and the extended nebula, thus making it a candidate double-envelope PN.
3.3.1 Light Curves
Figure 8 shows the ZTF and WISE light curves for CTSS 2 and K 3-5, with both objects showing very similar behavior. Over the past decade, the brightness of the objects in ZTF- and the redder bands have gradually increased by a factor of . An increase in brightness in ZTF- is also seen, albeit with much less prominence (flux increment by ). The light curve for K 3-5 has two additional intriguing features. First is the turn-over of the WISE light curves at around , unlike the optical light curve. Second, a significant bluing in the WISE bands (brightening of W1 compared to W2) is seen, contrary to the behavior in the optical. In CTSS 2, the WISE bands show no significant color dependence, unlike the ZTF light curves.
3.3.2 Follow-up Optical Spectroscopy
CTSS 2 and K 3-5 were observed with the Low-Resolution Spectrograph (LRS2) on the 10-m Hobby-Eberly Telescope (HET; Ramsey et al. 1998; Hill et al. 2021) located at the McDonald Observatory in West Texas, USA, as a part of a survey for faint PN central stars (undertaken by H.E.B, see Bond et al. 2023 and subsequent papers). CTSS 2 was observed four times: twice on 15th September, 2024 with 250 seconds exposure each and twice again on 15th October, 2024 with 300 seconds exposure each. K 3-5 was observed once on 10th August, 2024 with 900 seconds exposure.
The LRS2 instrument is described in detail by Chonis et al. (2014, 2016). Briefly, LRS2 provides integral-field-unit (IFU) spectroscopy with 280 -diameter lenslets that cover a field of view (FOV) on the sky. LRS2 is composed of two arms: blue (LRS2-B) and red (LRS2-R). All of our observations are made with the target placed in the LRS2-B FOV. The LRS2-B arm employs a dichroic beamsplitter to send light simultaneously into two spectrograph units: the “UV” channel (covering 3640–4645 Å with a resolving power of 1910), and the “Orange” channel (covering 4635–6950 Å with a resolving power of 1140). Data reduction was performed using the pipelines developed by G.R.Z.888https://github.com/grzeimann/Panacea,999https://github.com/grzeimann/LRS2Multi (Zeimann et al., in prep.).
The narrow-band (NB) images for both objects around the brightest emission lines in the spectrum: H and [O III] 4363, 4959, 5007 Å are shown in Figure 9. From the two-dimensional images, one-dimensional spectra were extracted for both the central core (nucleus) and the nebula. The corresponding apertures are shown within the NB images (white and black dashed apertures correspond to the nucleus and nebula, respectively). The extracted spectra for CTSS 2 and K 3-5 are shown in Figures 10 (top panel) and 11 respectively.
Both CTSS 2 and K 3-5 were also observed with DBSP on 13th May, 2024 with exposures of 15 minutes. It is different from the HET spectrum only for its wider coverage in the red arm. However, the HET spectrum is of higher quality and, being an IFU, enables more accurate analyses. Thus, we base our discussion on the HET spectrum in the following sections. We provide the DBSP spectrum separately in Appendix C.
3.3.3 Spectral Analysis: CTSS 2
CTSS 2 consists of a bright nucleus and a faint surrounding nebula. The left panel of Figure 9 shows the NB image of CTSS 2 around a few major emission lines. The bright core is very prominent and the faint nebula is also detected (especially in [O III] 4959, 5007 Å images). It is clear from these images that the nucleus is also very bright in the high-excitation emission lines, making it a double-envelope nebula.
We now turn to the one-dimensional spectra to understand the properties better (top panel in Figure 10). It is readily seen that the spectra (especially the relative strengths of the intensities of the major emission lines) of the nucleus and the nebula are markedly different. We now draw the attention of the readers to the strengths of H and [O III] 4363 Å relative to [O III] 4959, 5007 Å. In the nucleus, we find that the intensity of H is comparable to [O III] 5007 Å which is often attributed to very high electron densities. This is also supported by the high intensity of the [O III] 4363 Å emission line relative to the other [O III] lines.
To perform a more quantitative estimate, we define (as in Paper I) , which is related to the electron densities and temperatures (see Equation 3 in Paper I, taken from Osterbrock & Ferland 2006). Overall, a lower value of signifies very high densities and/or temperatures. We obtain for CTSS 2 nucleus. This value is strikingly small for a PN. With this value, any reasonable temperature would estimate electron densities of . Unlike the nucleus, the nebula as a much larger of . This is consistent with the expectation that the diffuse nebula should be significantly less dense than the core. However, this is still lower than in most PNe and indicates moderately high electron densities of .
We also find evidence of fast wind/shock activity in the nucleus. The top panel of Figure 12 shows a zoom-in of the base of the H emission lines for both the nuclear and the nebular spectra. Broad wings extending up to velocities of are seen in the nuclear spectrum. In contrast, the nebular spectrum is much narrower, showing that the wind activity is mostly localized to the core. But the wings of the nebular H still extend up to velocities of , which might indicate an extended influence of the nuclear activity. Note that such wind/shock may heat up the environment beyond that achievable through photoionization equilibrium, and contribute to lowering the inferred .
The continuum in the nuclear spectra contains absorption signatures (H and He), which we attribute to the central star. We also detect an emission doublet of C IV 5801, 5812 Å, typical of many hot stars. To infer the stellar parameters we performed a spectral fit. We computed a series of non-local thermodynamic equilibrium models of the type introduced by Werner et al. (2019) using the Tübingen Model-Atmosphere Package (Werner et al., 2003). They are plane-parallel and in radiative and hydrostatic equilibrium. They are composed of hydrogen and helium while carbon and nitrogen were included as trace elements keeping fixed the atmospheric structure. We assumed solar composition. Models were computed in the range = 50–90 kK and = 3.9–4.8. The best-fit parameters were determined by eye and found to be kK and (bottom three panels in Figure 10 shows the fit to the normalized spectrum). Using evolutionary tracks by Miller Bertolami (2016) this corresponds to stellar mass, radius, and luminosity of , , and . The corresponding age of the central star is inferred to be years. This makes it a candidate for one of the youngest post-AGB phase CSPN. We comment here on the possible “contamination” from any nebular continuum. We do not expect significant contamination from the faint diffused nebula, but any continuum associated with the central emitting core can potentially dilute the absorption lines, rendering the spectral fit suboptimal. As this emission is mostly unresolved, appropriate modelling of the emission lines is needed to correct for any such contamination. Such a detailed study is beyond the scope of this paper.
The Gaia parallax distance to CTSS 2 is kpc. The analysis of Bailer-Jones et al. (2021) yields a slightly different distance estimate of kpc. We use the inferred stellar parameters to get an independent estimate of the distance to CTSS 2. We first employ a simple manual comparison of the model to the observed spectrum, by varying the reddening and the distance. A good visual fit was obtained with (i.e. ) and kpc (see inset in the top panel of Figure 10). The distance is consistent with the Gaia distances.
The extinction value used above is consistent with that inferred from the Balmer line ratios. We first use the nebular spectrum to estimate . Following the same strategy as in Paper I, we get (the ratio of H to H) and which leads to extinction estimate of . For the nuclear spectrum, we obtain and which translates to . This is slightly higher than the nebular estimate but broadly consistent.
From the HST image in Stanghellini et al. (2016), we estimate the angular diameter of the nebula to be around . Assuming a distance to the nebula of kpc, gives the physical radius of the nebula is pc. This gives an average expansion speed of . This is marginally higher than the average expansion velocities of PNe (, Pereyra et al. 2013; Jacob et al. 2013).
3.3.4 Spectral Analysis: K 3-5
The right panel in Figure 9 shows the narrow band images on K 3-5. It is quite intriguing to observe that the nucleus is brighter than the nebula in Balmer and [O III] 4363 Å emissions, however, the reverse is true for [O III] 4959, 5007 Å images.
We proceed to perform a similar analysis as with CTSS 2 with the one-dimensional spectra (Figure 11): comparing the nuclear and nebular spectra and inferring the physical properties. The strengths of the emission lines are different between the nucleus and the nebula in a very similar way as in CTSS 2. The increased strength of the H and [O III] 4363 Å emission lines indicate a double envelope nature of the PN with a localized high-density emission core. This is in line with the inferences in Weidmann et al. (2018).
For the nebular spectrum, we get . This is at par with the typical values for evolved PNe. Additionally, [S II] 6716, 6731 Å lines are prominently detected in the spectrum. We obtain an intensity ratio of [S II] 6716/[S II] 6731. These two line ratios combined (from Osterbrock & Ferland 2006) yield electron densities of and temperature of kK. These properties are consistent with that of a PN.
For the nucleus, however, we derive a significantly lower of . For any temperature, this yields , significantly higher than the inferred nebular density, in line with it being a high-density emission core. We note here that weak [S II] lines are also detected in the nuclear spectrum, bearing a similar intensity ratio as the nebula. This is inconsistent with the low value. We thus conclude that this is just a contamination in the nuclear aperture from the overlapping nebula. Unlike CTSS 2, no absorption signature from the central star is detected in the nuclear spectrum preventing any parameter estimation of the CSPN. However, certain emission lines might arise from the central stellar system. For example, we detect C IV 5801, 5812 Å emission line in the nuclear spectrum (absent in the nebular spectrum). This doublet is often associated with either the photosphere of a hot CSPN (like CTSS 2) or an irradiated companion. If the CSPN is an emission line star, H or He emissions from the CSPN can contribute to the observed spectrum. However, further studies are needed to disentangle the nebular and the stellar origin.
Like CTSS 2, K 3-5 also has broad H emission wings in the nuclear spectrum, extending up to velocities of (bottom panel of Figure 12). On the other hand, the nebula profile is much narrower, it’s width probably being just limited by the instrument resolution. Note that faint [N II] 6548, 6583 Å are detected in both the nuclear and the nebular spectra (marked in the figure, though we suspect the former to be just nebular contamination like the [S II] lines). In the same figure, we see an emission signature at in the nuclear spectrum. This can be the He II 6527 Å line corresponding to the to transition (see for example Lee et al. 2001).
We infer a very high reddening for this object. We infer and from the nuclear spectrum. The nebular spectrum also yields consistent values of and respectively. This translates to . This explains the apparent red continuum of the nuclear spectrum. We note that such a high reddening revises the inferred values of for the nucleus and the nebula to and . None of our conclusion changes because of this, apart from increasing the density estimation for the nucleus to .
3.3.5 Spectral Energy Distribution: Infrared Excess
To look for the presence of dust in these systems, we investigate their spectral energy distributions (SED). We retrieved the SED of CTSS 2 and K 3-5 (around ′′ from the target coordinate) from the VizierR online photometry viewer tool. 101010http://vizier.cds.unistra.fr/vizier/sed/ The result is presented in Figure 13. Some scatter is evident in the data, especially for K 3-5, which is possibly due to the extended nature of the objects leading to suboptimal photometry. Nevertheless, the presence of significant infrared excess in both the objects is evident.
For a more quantitative estimate, we perform multi-component blackbody fit to the data points with wavelengths . For CTSS 2, a reasonable fit was obtained with a two-component fit. The stellar component was subtracted and distance of kpc was assumed. For K 3-5, we assumed the Gaia eDR3 parallax distance of pc, and a reasonable fit was only observed with three components. The fit parameters of the different components are presented are annotations in the same Figure 13.
We interpret the various blackbody components as representative of extended dust structures. We find that both CTSS 2 and K 3-5 appear to have warm dust with temperature in the range of K and radius . We also find evidence of extended colder dust features. For K 3-5 it appears to be extended to around a thousand AU, with K. For CTSS 2, the second component appears hotter and more localized to around . But this might be due to inadequate coverage in the redder wavelengths, thus missing out on the more extended cooler dust components.
3.3.6 Discussion
The observed photometric evolution can result from several different possible scenarios. Firstly, the observed light curve can be a part of a longer periodic variability due to binarity. Given we do not see the turnover in the light curve over the current baseline (except possibly in WISE data for PN K 3-5), the lower bound to the period is years. The observed variability, thus, would have to be from the irradiation effect. However, such a large amplitude variation is unlikely at such a wide orbit, making this scenario less probable.
Motivated by the inferred stellar parameters of CTSS 2, we consider a rare but very important scenario: late thermal pulses (LTP). LTP designates the occurrence of a He flash in the post-AGB phase, but prior to entering the white dwarf cooling track. The subsequent evolution depends strongly on the conditions at the He flash. But, a typical post-LTP evolution involves expansion of the star (and thus increase in luminosity) along with decrease in temperature as it re-enters the AGB phase (see for example the theoretical tracks discussed in Lawlor 2023). A past example includes FG Sge, which showed extensive photometric brightening (before it turned into an R Coronae Borealis star), at a very similar rate as CTSS 2 (see, for example, Figure 3 in van Genderen & Gautschy 1995). The cooling of the star can also explain the strong reddening observed in the light curves of both CTSS 2 and K 3-5. Note that an LTP model of a 0.625 star was used to explain the evolution of FG Sge (Jeffery & Schönberner, 2006), which agrees very well with the inferred mass of the CSPN of CTSS 2. Additionally, an LTP candidate with parameters very close to that of CTSS 2 was reported in Werner et al. (2025, GSC08265). However, unlike CTSS 2, the spectrum of GSC08265 shows extensive signatures of C and O, indicative of the dredge-up. But, LTP models indicate that the dredge-up occurs only much later (post or near attainment of the minimum temperature, Jeffery & Schönberner 2006). In most cases of LTP (like LTP type III, Lawlor 2023), the photospheric composition remains nearly unchanged during the initial post-LTP phases of the star. CTSS 2 stands as a candidate to an LTP in this stage. K 3-5 might also be a LTP candidate, given the similarity in light curve behavior. But without any spectroscopic identification of the CSPN, the inference is more uncertain.
We now explore other possible scenarios. One object showing a similar long-term photometric evolution is IC 4997. It shows decades-long photometric evolution comprising of both dimming and brightening episodes (Aller & Liller, 1966; Kostyakova & Arkhipova, 2009; Arkhipova et al., 2020). It is a young planetary nebula. The CSPN has been classified to be a possible weak emission-line star (WELS) primarily owing to the presence of C IV 5810, 5812 Å and C III 4650 Å emission lines (note that the former is present in both CTSS 2 and K 3-5). The current understanding attributes the observed photometric behavior to changing wind/mass-loss episodes. As discussed, both objects show evidence of wind through broad H lines. Additionally, significant photometric reddening during brightening has been observed in IC 4997 (Arkhipova et al., 2020), similar to both these objects. Alongside, IC 4997 also shows spectroscopic evolution with changing line ratios indicating changing and over time, also attributed to wind and shocks. This warrants a careful comparison of the current and the archival spectra111111Comparison of the current and past spectra indeed indicates spectroscopic evolution of both CTSS 2 and K 3-5. The spectra of both of them on HASH (and in Stanghellini et al. (2016) for CTSS 2) has H strength much less than the neighboring [O III] lines, and a weak [O III] 4363 Å. This might indicate an increase in associated with the brightening. But we do not establish it with confidence because of a lack of knowledge about the details of the archival spectra (trace selection, nucleus vs nebula).. Because of these reasons, IC 4997 also has a very dense nebula, similar to the core of both these objects.
Long-term dust activity can also lead to photometric evolution. A specific example is M 2-29 (Hajduk et al., 2008; Miszalski et al., 2011b). Both CTSS 2 and K 3-5 bear many similarities to M 2-29: broad H wings indicating wind activity, high-density emission-line core, and presence of emission lines from the CSPN. The presence of extensive dust structures in both objects is clearly evident from the SED. Brightening in such cases results from recovery from heavy dust obscuration. However, we expect blue-ing instead of reddening in such cases. Thus, this scenario appears unlikely though longer photometric monitoring is desired.
The high-density emission core in these nebulae is itself an intriguing aspect. This is the property of the so called EGB 6-type PNe (Liebert et al., 2013) bearing Compact Emission Knots (CEKs), hypothesized to be an irradiated dense accreted envelope around a wide-orbit companion to the CSPN white dwarf. Thus, this may be an indirect indication of the binarity of the CSPNe. For K 3-5, the bipolar nature of the nebula also supports multiplicity of the CSPN. Another explanation involving a companion might be a symbiotic interaction, where such high density cores are common. The lack of the spectral signatures of a Mira giant in the red end (see the DBSP spectrum with a better red coverage, Figure C1), however, speaks against them being the more common D-type symbiotic system. They can be the less common S or D’-type symbiotics. Presence of these systems inside PNe are rare. Other confirmatory signatures of symbiotic behavior, like the Raman-scatter O VI emissions, are also not evident. Further multi-wavelength spectroscopic observations are required to understand these objects.
3.4 LoTr 1
LoTr 1 is a True PN in HASH and has a diameter of 145′′. It’s CSPN, AB Lep, is a known variable.
3.4.1 Light Curves
Long-term photometric modulation was reported in Martínez et al. (2022), with a period of 11.28 years. They present LoTr 1 as a RS CVn system: a close binary with a giant companion where the magnetic activity cycle of the giant star leads to long-period photometric variability. On short timescales, it shows a period of 6.4 days, which has been attributed to the rotation of the giant star (Tyndall et al., 2013).
The ZTF light curve shows a long-lasting ‘dip’ (left panel in Figure 14). This is likely a segment around the minimum of the longer periodic variability (the time of the minima and the amplitude are consistent with that expected from Martínez et al. 2022). We successfully recover the signal in the season-wise de-trended ZTF light curve (the right two panels in the same figure) as the dominant periodogram peak. Note that C25 also detected this period, but, unlike this work, it did not appear as their dominant peak (their Figure 14).
We now note that LoTr 1 is similar to LoTr 5, which consists of an unresolved hot (pre-)white dwarf and rapidly rotating magnetically active G5 giant pair. For the latter system Kővári et al. (2019) uncovered a long-term photometric variability which appears in phase with the orbital radial velocity of the binary ( yr). Thus, one can also speculate that the long-term variability in LoTr 1 might also reflect the orbital period of the binary system. Thus, unlike Martínez et al. (2022), we propose that this is a wide binary with a rapidly rotating giant companion.
3.4.2 HST/COS Spectroscopy and Spectral Energy Distribution
The central star of LoTr 1 was observed for 2299 s on March 29, 2023, with the Cosmic Origins Spectrograph (COS) aboard the Hubble Space Telescope (HST), using the G130M grating centered on 1291Å (dataset LEZCI2020, PI: N. Reindl). Another 1640 s exposure was taken with HST/COS on July 30, 2023, using the G140L grating () that covers Å (dataset LEZCI5010, PI: N. Reindl). The spectra were retrieved from the MAST archive.
The hot central star of LoTr 1 turns out to be a spectroscopic twin of the extremely hot white dwarf in UCAC2 46706450 (Werner et al., 2020) as it becomes obvious when looking at Fig. 15 where we compare the UV spectra of both stars. The lack of observed Fe V lines as well as absence of the C III multiplet at 1175 Å imposes a lower limit of 100 000 K (Werner et al., 2020). In the G140L spectrum a weak He II Å can be detected which suggests that the central star of LoTr 1 is either a very hot DAO white dwarf or a O(H)-type pre-white dwarf. Interestingly, the companion of UCAC2 46706450 is also a rapidly rotating giant star.
In order to estimate the radii of both stars we carried out a two component fit to the spectral energy distribution (SED) of LoTr 1, employing the SED fitting routine described in Heber et al. (2018) and Irrgang et al. (2021). For the cool star we used PHOENIX models calculated by Husser et al. (2013) and for the hot central star the metal free model grid computed by Reindl et al. (2023). We fixed the of the central star to 105 000 K (i.e. to the value that was derived for UCAC2 46706450 by Werner et al. 2020), and then let the let the angular diameter, reddening, surface ratio of both stars, as well as the of the cool companion vary freely. By that we derive a reddening of mag121212Fitzpatrick et al. (2019) employs , which is the monochromatic equivalent of usual , using the wavelengths 4400 Å and 5500 Å, respectively. For high effective temperatures such as for the stars in our sample is identical to . and a radius of 0.039 for the hot central star (identical to UCAC2 46706450), indicating that it is just about to enter the white dwarf cooling sequence. For the cool giant star we derive K and a radius of 9.7 . These parameters are broadly consistent with a subgiant star of mass in the range of ,131313We compared the stellar parameters to Modules for Experiments in Stellar Astrophysics (MESA, Paxton et al. 2011, 2013, 2015) evolutionary tracks queried from MESA Isochrones and Stellar Tracks (MIST, Dotter 2016; Choi et al. 2016). Owing to the preliminary nature of the analysis we do not show the comparison. This will be presented elsewhere with spectroscopic observations. which is also very similar to the companion of UCAC2 46706450. With the 6.48 day rotational period, this implies a surface velocity of .
3.4.3 Discussion: Growing Evidence of Interaction in Wide Binaries during AGB?
We briefly discuss here a growing class of objects of which LoTr 1 is a likely members: a hot white dwarf with a rapidly rotating giant companion at a wide orbit. In PN literature, such systems are known as Abell 35 type CSPNe (after the first such system, Jacoby 1981; Jasniewicz & Acker 1988). There are several such CSPNe known (see Bond & Zeimann 2024 and their Section 8 for a concise account of such CSPNe). We note here that WeSb 1, the recently discovered CSPN with orbiting dust and debris disk (Bhattacharjee et al., 2025; Budaj et al., 2025) is also likely to be a similar system. Recent high resolution optical spectrum has shown that the subgiant companion of WeSb 1 is rapidly rotating with (details to be presented elsewhere, Bhattacharjee et al. in preparation). A likely formation channel of these systems is mass transfer from the white dwarf progenitor AGB to the companion which spins it up (see, for example, Jeffries & Stevens 1996). These systems, thus, potentially hold valuable information about binary interactions at wide orbits.
The occurrence of such system will be less likely down the white dwarf cooling track as the companion is expected to eventually spin down. The discovery of UCAC2 46706450, however, shows that such systems can be found even after the associated PN has faded away. Thus, we speculate that a population of such sources exist outside of PNe and are yet to be found and studied.
3.5 The other long-timescale variables
In this section, we briefly introduce the remaining five long-timescale variables. Further spectroscopic and photometric studies of these objects are underway and will be presented in future works. The ZTF light curves are shown in Figure 17. Except for Kn J1857.7+3931 (henceforth, Kn J1857) and IPHAS J192717.94+081429 (henceforth, IPHAS J1927), all of the long-timescale variables are True PN on HASH with resolved nebulae. Kn J1857 has now been reclassified to ‘object of unknown nature’ on HASH and IPHAS J1927 is a Possible PN. Neither of the two has a resolved nebula (thus either very young PNe or some stellar PN mimic).
We first discuss the True PNe. PN K 3-20 shows a brightening event, with a larger amplitude in the ZTF band. Tan 2 shows a singular eclipse-like dip in the band. Unfortunately, the band data for this object did not pass the photometric quality cuts. The light curve of WHTZ 3, albeit noisier, shows low amplitude long-term modulation. The possible causes for variability of these objects are uncertain. The observed light curves can be a part of even longer timescale periodic variability associated with binarity. For example, the behavior in PN K 3-20 may result from irradiation effect in a long-period wide binary (similar to NGC 6833). Further long-term photometric monitoring is needed to test these scenarios. Concerning the remaining two objects, Kn J1857 shows a gradual dimming, by in flux, with similar amplitude in both the ZTF bands. The variability in IPHAS J1927, on the other hand, is more irregular. Similar kinds of photometric behaviors are often seen in symbiotic systems or cataclysmic variables (CVs). Being only a ‘Possible PN’, spectroscopic data is necessary to determine the true nature of this object.
4 Summary and Conclusions
This is the second of a series of two survey papers undertaking a systematic study of the variability of the central stars of planetary nebulae (CSPNe) using the optical light curves from the Zwicky Transient Facility (ZTF). In Paper I, we applied appropriate variability metrics to identify significantly variable sources. This paper focuses on 11 objects among them, which show long-timescale variability. For six of these objects, we present follow-up spectroscopic/photometric observations and provide extensive discussion, which is expected to serve as the first step in understanding the objects. We also present the remaining six objects and encourage follow-up studies by the community.
The first object we discuss is NGC 6833 (Section 3.1). It shows ‘triangle-shaped’ brightening in the ZTF , , and WISE W1 and W2 bands, occurring at a period of days. The more intriguing aspect is the anticorrelated behavior in the ZTF -band, which shows shallow dips nearly at the same epochs as the peaks in the redder bands. We speculate that this might be a long-period eccentric binary, with the photometric period corresponding to the orbital period. This makes it the candidate for the fourth to the longest orbital period inside a PN. Though the irradiation effect can possibly explain the brightenings, the simultaneous dips in -band is puzzling. We speculate this might be due to dust production through certain stellar interactions at orbital periastron.
Long-periodic modulation was also detected in the light curves of NGC 6905 and Kn 26 (Section 3.2. The former showed a significant peak at days and the latter at days. Secondary significant peaks were also detected at days and days were also detected. The modulation appears broadly sinusoidal. This makes both the objects candidates for long-period binary inside CSPNe. Their periods (as well as that of NGC 6833) are short enough for them to be post-CE systems, making them candidates for post-CE wide binaries. Kn 26 displays an additional periodicity at 1.18 hr, which we verify through follow-up photometry. The origin of this period is uncertain. It can be of close binary (in which case the interpretation of the long period as a wide orbit binary makes it a candidate triple system) or pulsational origin.
We present CTSS 2 and PN K 3-5, which show slow and long-term brightening (Section 3.3). The objects brightened by almost a factor of in the ZTF -band and WISE bands, but much less so (factor of ) in the -band, thus undergoing significant reddening. Spectroscopy of both the nucleus and the nebula shows that both objects host high-density () emission-line cores. Thus they are EGB 6-type PNe where accreted ejecta on a wide companion is irradiated, or symbiotic systems. The optical spectrum of CTSS 2 displays stellar signatures. A spectral fit yields a temperature of kK and a stellar mass of . This translates to a post-AGB age of years, making it one of the youngest CSPN known. Unfortunately, the spectrum of K 3-5 does not show adequate stellar features preventing any spectroscopic analysis of the pre-white dwarf. We speculate that these might be late thermal pulse events. In this case, the light curves suggest that the stars are cooling down and expanding, moving back to another AGB phase. This is a very rare occurrence and would indicate the success of long-baseline surveys like ZTF to discover these objects. Other possible explanations include wind/dust activity of the CSPN system.
We present new results on LoTr 1 (Section 3.4), which is a known variable showing a long-period modulation at years and a shorter period at days. The latter is attributed to the rotational modulation of a fast-rotating cool giant companion. The longer period might be the orbital period. Ultraviolet spectroscopy with HST shows that the hot component in the system is a spectroscopic twin of the extremely hot white dwarf in UCAC2 46706450, with a temperature of kK. This hints at the presence of a whole class of objects with fast-rotating giant companions to very hot white dwarfs. Further studies are underway and will be presented elsewhere. A fit to the SED reveals the stellar parameters of the giant to be K and radius of .
Finally, we briefly discuss the observed variability in the remaining five objects (Section 3.5). They display a variety of photometric behaviors: long timescale dimming, single brightening event, single dip event, or more erratic variability. Follow-up studies are necessary to understand these systems.
Overall, this study shows the effectiveness of long-baseline surveys in uncovering the variety of photometric behaviors observed CSPNe. Efforts need to be made to not only understand the objects individually through follow-up observations but, if possible, to establish evolution links to different classes of objects. For example, the peculiar photometric behavior of NGC 6833 may shed significant light on the formation and evolution of post-CE wide binary systems. Note here that ZTF can only access the north sky. The majority of the PNe are in the southern sky, near the galactic center. Upcoming surveys like LSST will observe these targets, and we expect to find several more such intriguing systems. Simple variability metrics and metric spaces (as demonstrated in Appendix A) will significantly boost such discoveries.
5 Acknowledgements
This work is based on observations obtained with the Samuel Oschin Telescope 48-inch and the 60-inch Telescope at the Palomar Observatory as part of the Zwicky Transient Facility project. ZTF is supported by the National Science Foundation under Grants No. AST-1440341 and AST-2034437 and a collaboration including current partners Caltech, IPAC, the Oskar Klein Center at Stockholm University, the University of Maryland, University of California, Berkeley, the University of Wisconsin at Milwaukee, University of Warwick, Ruhr University Bochum, Cornell University, Northwestern University, and Drexel University. Operations are conducted by COO, IPAC, and UW.
This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC; https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular, the institutions participating in the Gaia Multilateral Agreement.
We are grateful to the staffs of Palomar Observatory and the Hobby-Eberly Telescope for assistance with the observations and data management. The Liverpool Telescope is operated on the island of La Palma by Liverpool John Moores University in the Spanish Observatorio del Roque de los Muchachos of the Instituto de Astrofisica de Canarias with financial support from the UK Science and Technology Facilities Council.
The Low-Resolution Spectrograph 2 (LRS2) on HET was developed and funded by the University of Texas at Austin McDonald Observatory and Department of Astronomy, and by Pennsylvania State University. We thank the Leibniz-Institut für Astrophysik Potsdam (AIP) and the Institut für Astrophysik Göttingen (IAG) for their contributions to the construction of the integral field units. We acknowledge the Texas Advanced Computing Center (TACC) at The University of Texas at Austin for providing high performance computing, visualization, and storage resources that have contributed to the results reported within this paper.
The Isaac Newton Telescope is operated on the island of La Palma by the Isaac Newton Group of Telescopes in the Spanish Observatorio del Roque de los Muchachos of the Instituto de Astrofísica de Canarias
S.B. thanks Frank J. Masci and Zachary P. Vanderbosch for useful discussions and suggestions regarding solving the issues with ZTF forced photometry on extended sources. S.B. also thanks Jim Fuller, Charles C. Steidel, Lynne Hillenbrand, and Adolfo Carvalho for useful discussions on methods and science. S.B. acknowledges the financial support from the Wallace L. W. Sargent Graduate Fellowship during the first year of his graduate studies at Caltech. N.C. was supported through the Cancer Research UK grant A24042.
DJ acknowledges support from the Agencia Estatal de Investigación del Ministerio de Ciencia, Innovación y Universidades (MICIU/AEI) under grant “Nebulosas planetarias como clave para comprender la evolución de estrellas binarias” and the European Regional Development Fund (ERDF) with reference PID-2022-136653NA-I00 (DOI:10.13039/501100011033). DJ also acknowledges support from the Agencia Estatal de Investigación del Ministerio de Ciencia, Innovación y Universidades (MICIU/AEI) under grant “Revolucionando el conocimiento de la evolución de estrellas poco masivas” and the the European Union NextGenerationEU/PRTR with reference CNS2023-143910 (DOI:10.13039/501100011033).
Appendix A The Von-Neumann and Pearson Skew metric space
In Appendix B of Paper I, we presented a two-dimensional metric space defined by the variability metrics von Neumann statistics and Pearson Skew (vonN and SkewP, equations B1 and B2 in Paper I respectively). We demonstrated the ability of the metric space to identify ‘uniquely’ variable light curves, like WeSb 1, from a larger pool of objects. We briefly review the metric space and its functionality. The Vonn metric quantifies the randomness in the time-series data. Theoretically, an adequately sampled completely random data should have (von Neumann, 1941). SkewP, on the other hand, quantifies the directionality of the variability ( signifies dips/eclipses, and indicates outbursts). Non-variable sources with just photon shot noise, thus, cluster around (slightly lower at in practice) and .
In Paper I, we discussed the position of the short-timescale variables in this metric space. We now focus on the long-timescale variables. Figure A1 is the same as Figure B1 in Paper I with all the 11 long-timescale variables marked individually. The notable feature is that almost all the sources (except WHTZ 3, where the light curve is noisier) have low values of the Vonn metric (). This is because the variability in these sources is well resolved in ZTF cadence and, thus, non-random. Based on the light curves presented in the main paper, the values of SkewP are easy to interpret. Sources showing brightening events, like NGC 6833, CTSS 2, PN K 3-5, and PN K 3-20, have . Whereas objects showing dips like AB Lep or Tan 2 have . The variability in NGC 6905 and Kn 26 is symmetric, thus they don’t have significant positive or negative SkewP values. The SkewP values of the more complex variables like PN G044.3+04.1 can be difficult to interpret. Nevertheless, most of the long-timescale variables stand out from the other objects in the space and are easily identifiable. This thus potentially serves as a great tool to filter interesting sources from large data sets of light curves.
Appendix B NGC 6905 and Kn 26 Periodicity Tests
We briefly describe here the various tests performed to confirm the reality of the inferred periods of NGC 6905 and Kn 26. Though not shown, the periods were recovered from standard photometry data too, ruling out any difference image artifact. To check if the observed variability is due to seasonal variation in atmospheric conditions, we compared the atmospheric seeing to the observed flux. No correlation was found. To test if the periods arise from the seasonal breaks or yearly cycles in the survey, we randomized the flux values keeping the time stamps constant, and manually investigated the periodogram in several iterations. No significant peak was obtained and the periodograms were as good as noise. This is also reflected as a negligible false alarm probability (FAP) calculated using the inbuilt bootstrap method in astropy.timeseries.
We then query ZTF standard photometry light curves of all Gaia sources within a couple of arcminutes from the targets ( objects for each source). Visually, none of the neighboring objects showed the variability seen in the target sources. These include sources behind the extended nebulae of the PNe, thus ruling out any spurious nebular activity resulting in the observed modulation. We applied the same period search procedure on all neighboring objects’ light curves and manually inspected the periodograms. For the most significant period in each of the resultant periodograms, we investigate the distribution of the FAPs. None of the objects showed the periods inferred in our targets with any reasonable FAP or periodogram signal to noise. The only two spurious periods with low FAP in some of the objects are one year and days (the ZTF baseline), but nowhere close to the inferred periods in the two targets.
For a final test, we queried the difference-image forced photometry data from the Asteroid Terrestrial-impact Last Alarm System (ATLAS) survey141414We queried both the difference and science-image forced photometry data from the online service at https://fallingstar-data.com/forcedphot/ (Shingles et al., 2021). Owing to their warnings about the latter, we primarily use the former but use the science-image data for estimation of the reference flux and normalization. We reject all data where err. for both the targets and perform period search. The results are presented in Figure B1. Qualitatively, similar variability as in ZTF is recovered in ATLAS. This rules out with confidence both any ZTF-specific systematics and region-specific atmospheric seeing variation inducing the variability. Significant peak in the periodogram appears for NGC 6905, albeit at a slightly higher period of days. The disagreement may result from the ATLAS data being significantly noisier than ZTF. A secondary peak at a similar period as in ZTF is recovered. For Kn 26, we successfully recovered the dominant ZTF period of days. The secondary peak in ATLAS has a lower signal-to-noise, but the corresponding period is broadly in agreement with that in ZTF. We briefly note here that ATLAS data is too noisy to detect the short period in Kn 26.
Appendix C DBSP Spectra of CTSS 2 and PN K 3-5
Figure C1 shows the DBSP spectrum for CTSS 2 and PN K 3-5. The same reduction was procedure as with WeSb 1 in Paper I and NGC 6833 was followed. With a slit spectrograph, it is difficult to separately identify the nucleus and the nebula. Thus, we choose a fiducial trace containing the main emission core and extract the corresponding one-dimensional spectra. The only advantage over the previously presented HET spectrum is its better spectral coverage in the red. This enables us to check for obvious indication of a giant star, which may indicate a symbiotic system. Such signatures are not present in the spectrum.